OpenStudy (anonymous):

Please help i have no idea how to do this at all :'( Find the quadratic function y=ax^2+c with a graph that includes (0,1/3) and (-3,82/3) A. y=3x^2 + 1/3 B. y= -3x^2 + 2/3 C. y= 3x^2 - 1/3 D. none of these

OpenStudy (***[isuru]***):

plug the given coordinates to the equation u got considering ( 0 . 1/3 ) y = ax^2 + c 1/3 = a(0)^2 + c so c = 1/3 considering (-3 , 82/3) y = ax^2 + c 82/3 = a (-3)^2 + 1/3 82 = 27a + 1 27a = 81 a = 3 so.. the equation is y= 3x^2 + 1/3 got it ?

OpenStudy (anonymous):

oh wow its that simple? lol i feel so dumb thank you SO much!!!!!!

OpenStudy (***[isuru]***):

u r welcome...